# How to Bridge Structural and Temporal Heterogeneity in Link Prediction? A Contrastive Method

### v1

@@ 5.3. Contrastive Learning Approaches
Contrastive learning (Chen et al., 2020; Liu et al., 2023; Wu et al., 2023b) represents a promising avenue for link prediction by leveraging unsupervised representation learning with positive and negative sample pairs. Notable work like GraphCL (You et al., 2020) employs random graph augmentations to generate positive and negative pairs of graph instances, facilitating better representation learning. HeCo (Wang et al., 2021a) designs a novel co-contrastive mechanism that effectively preserve both meta-path information and network schema information. LGCL (Zhang et al., 2023b) converts the link prediction into node classification by transforming subgraphs into line graphs, and applying contrastive learning to maximize mutual information between the original subgraph and its corresponding line graph. Focus on incomplete graphs, SMiLE (Peng et al., 2022) employs a dual-view encoder to capture the structural and contextual information of entities. Positive samples are generated based on context subgraphs, while negative samples are derived from both intra-schema and inter-schema contexts. Instead of emphasizing on embedding individual nodes, Xu et al. (Xu et al., 2023) address a task of discriminating between groups based on structural similarity, and optimize their embeddings to differentiate these groups. While these graph contrastive learning approaches enhance the discriminative power of link prediction models, they generally adopt a coarse-grained perspective and do not address fine-grained nonuniformity.
@@ 6. Conclusion
In this paper, we propose a novel contrastive learning-based link prediction method, CLP, which develops a multi-view hierarchical self-supervised architecture to effectively characterize spatial and temporal heterogeneity. First, for spatial heterogeneity, we introduce a structural feature modeling component to identify topological distribution patterns at node- and edge- level. Next, for temporal heterogeneity, we devise a dual-channel architecture that represents the evolutionary progressions within dynamic graph topologies over both the long and short terms. Lastly, we design a multi-view hierarchical representation architecture to integrate the heterogeneity paradigms, effectively encoding spatial and temporal distribution complexity. The efficacy of our model for temporal heterogeneous link prediction is validated through experiments on four public datasets, demonstrating its significantly superiority over state-of-the-art methods in link prediction tasks. Our future research aims to extend the application of THNs to diverse domains for advanced network analysis and prediction. Particularly, we plan to incorporate multimedia data, such as images and videos, into THNs to enrich our exploration and comprehension of multimedia content within these networks. By utilizing the dynamic features of THNs, we can delve into complex pattern recognition and anomaly detection in multimedia streams.
@@ Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China (2020YFB1406902), the Key-Area Research and Development Program of Guangdong Province (2020B0101360001), the Shenzhen Science and Technology Research and Development Foundation (JCYJ20190806143418198), the Major Key Project of PCL (Grant No. PCL2021A02), and the Fundamental Research Funds for the Central Universities (HIT.OCEF.2021007). Professor Hui He is the corresponding author (email: hehui@hit.
edu.cn).

### v3

@@ 5.3. Contrastive Learning Approaches
Contrastive learning (chen2020simple; LiuJPZZXY23; WuLTGL23) is a promising approach for link prediction, utilizing unsupervised representation learning with positive and negative sample pairs. Notable work like GraphCL (you2020graph) employs random graph augmentations to generate pairs of graph instances for improved representation learning. HeCo (WangLHS21) introduces a co-contrastive mechanism preserving both meta-path and network schema information effectively. LGCL (zhang2023line) transforms link prediction into node classification by using line graphs and applying contrastive learning to maximize mutual information. SMiLE (PengLXXWP22) focuses on incomplete graphs by employing a dual-view encoder to capture structural and contextual information for entity representation. Studies in (XuDXJ23) differentiate between groups based on structural similarity rather than individual nodes, optimizing their embeddings accordingly. Nevertheless, these methods enhance link prediction models’ discriminative power but often overlook fine-grained nonuniformity.
@@ 6. Conclusion
In this paper, we propose a novel contrastive learning-based link prediction method, CLP, which develops a multi-view hierarchical self-supervised architecture to effectively characterize spatial and temporal heterogeneity. First, for spatial heterogeneity, we introduce a spatial feature modeling component to identify topological distribution patterns at node- and edge- level. Next, for temporal heterogeneity, we devise a dual-channel architecture that represents the evolutionary progressions within dynamic graph topologies over both the long and short terms. Lastly, we design a multi-view hierarchical representation architecture to integrate the heterogeneity paradigms, effectively encoding spatial and temporal distribution complexity. The efficacy of our model for temporal heterogeneous link prediction is validated through experiments on four public datasets, demonstrating its significantly superiority over state-of-the-art methods in link prediction tasks. Our future research aims to extend the application of THNs to diverse domains for advanced network analysis and prediction. Particularly, we plan to incorporate multimedia data, such as images and videos, into THNs to enrich our exploration and comprehension of multimedia content within these networks. By utilizing the dynamic features of THNs, we can delve into complex pattern recognition and anomaly detection in multimedia streams.
@@ Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China (2020YFB1406902), the Key-Area Research and Development Program of Guangdong Province (2020B0101360001), the Shenzhen Science and Technology Research and Development Foundation (JCYJ20190806143418198), the Major Key Project of PCL (Grant No. PCL2021A02), and the Fundamental Research Funds for the Central Universities (HIT.OCEF.2021007). Professor Hui He (hehui@hit.edu.cn) and Professor Weizhe Zhang (wzzhang@hit.edu.cn) are the corresponding authors.
Appendix
Appendix A Algorithm
The detailed algorithm of our CLP is presented in Algorithm 1.
Algorithm 1 Link Prediction with CLP.
0: Final node representation embedding: $\mathbf{u}_{a}$.
1:for Each train iteration do
2:  Parameters initialization.
3:  for $\mathcal{G}^{t}$ in $\mathcal{G}$ do
4:   Construct subgraph $\mathcal{G}^{rt}$ based on the edge type $r\in R$;
5:   for Every subgraph $\mathcal{G}^{rt}$ associated with edge type $r$ do
9:    Calculate the node-level loss $\mathcal{L}_{Node}^{+}$ and $\mathcal{L}_{Node}^{-}$ via Eq. (5);
10:   end for
11:   Calculate the edge-level weight $\delta_{a}^{rt}$ for any node $a$ via Eq. (6);
14:   Calculate the edge-level loss $\mathcal{L}_{Edge}^{+}$ and $\mathcal{L}_{Edge}^{-}$ via Eq. (9);
15:  end for
18:  Calculate the time-level loss $\mathcal{L}_{Time}^{L+}$ and $\mathcal{L}_{Time}^{S+}$ via Eq. (12);
19:  Calculate the total loss $\mathcal{L}_{total}$ via Eq. (13) and Eq. (14).
20:end for
21:return $\mathbf{u}_{a}$
Table 3. Comparisons of Baseline Models and Our Proposed Approaches.
Model Key Components
Static Homogeneous SEAL (zhang2018link) Node Labeling, Subgraph Extraction, GNN Graph Learning
Baselines VGNAE (ahn2021variational) Graph Autoencoder, Variational Inference, GCN Graph Learning
Static Heterogeneous Metapath2Vec (dong2017metapath2vec) Metapath-guided Random Walks, Skip-gram Modeling
GATNE (cen2019representation) Attributed Multiplex Networks, GCN Graph Learning
Dynamic Homogeneous TGAT (XuRKKA20) Graph Attention Network, Self-attention Mechanism
TDGNN (qu2020continuous) GCN Graph Learning, Temporal Aggregator, Edge Aggregator
Dynamic Heterogeneous THAN (li2023memory) Dynamic Transfer Matrix, Self-attention Mechanism
THGAT (zhang2023dynamic) Neighborhood Type Modeling and Aggregation, Temporal Dynamics Integration
Ours CLP Node-level Feature Modeling, Edge-level Feature Modeling, Temporal Information Modeling
CLP-N Edge-level Feature Modeling, Temporal Information Modeling
CLP-E Node-level Feature Modeling, Temporal Information Modeling
CLP-T Node-level Feature Modeling, Edge-level Feature Modeling
Appendix B More Detailed Settings
B.1. Implemention Details
In this section, We provide more detailed fine-tuning of our hyper-parameters and baseline configurations. Specifically, the dimension $d$ of node embeddings is explored within the range of $\left\{8,16,32,64,128\right\}$. Moreover, we search for the optimal values of the balance coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ in the range of $1e\left\{-1,-2,-3,-4,-5\right\}$; the number of attention heads $h$ in the range of $\left\{1,2,4,8,16\right\}$; and the temparature coefficient $\tau$ in the closed interval of $[0.04,0.12]$ with a step size of $0.02$. By default, after fine-tuning, we adopt the following hyperparameter settings, wherein the optimal values of $d,\lambda_{\cdot},h,\text{ and }\tau$ are set to $32,1e-3,4,\text{ and }0.1$, respectively.
B.2. Datasets and Metrics
This section will present the details and processing of the datasets, along with the evaluation metrics. The dataset statistics are presented in Table 4.
B.2.1. Dataset Description
  • Math-overflow [^1] http://snap.stanford.edu/data/sx-mathoverflow.html.: Math-overflow serves as an interactive website for mathematics enthusiasts. It adopts a forum-style model for communication and interaction, fostering an online community of expert mathematicians. This dataset comprises 2350 days of user interactions, which can be categorized into three types: question and answer exchanges, question and comment discussions, and answer and comment engagements. For analytical purposes in our experiments, we divide this dataset into 11 snapshots using a time window of 124 days.
  • Taobao (DuWYZT19): Taobao is characterized by three types of node: user (U), product (I), and topic (T), along with three types of links: U-I, U-T, and I-T, which signify users’ interactions on cloud-themed products in the Taobao App from April 1 to May 31, 2008. For our experimentation, this dataset is divided into five snapshots using a time window of 12 days.
  • OGBN-MAG (hu2020open): OGBN-MAG is a subset of Microsoft Academic Graph (MAG), which contains four node types (papers, authors, institutions, and research areas) and four types of relationships (authors affiliated with institutions, authors writing papers, papers citing other papers, and papers associated with research areas as their topics). In our experiments, a THN is extracted from OGBN-MAG for the period from January 1st to 10th, 2010, divided into time slots with 1000 edges of each type to create the dataset for the experiment.
  • COVID-19222https://coronavirus.1point3acres.com/en: COVID-19, sourced from 1point3acres, provides state- and county-level daily case reports, including confirmed cases, new cases, deaths, and recovered cases. We select the daily new COVID-19 cases as our THN data. This dataset consists of two node types: (state and county) and three relationships: (state includes county, state adjacency, county adjacency). In our experiments, the constructed THN spans from May 1 to 21, 2020, with 21 time snapshots, limiting each relationship type to a maximum of 2000 edges per snapshot.
Table 4. Statistics of Datasets.
Dataset Math-overflow Taobao OGBN-MAG COVID-19
# Nodes 24818 29475 17269 2165
# Edges 506550 63367 40000 131649
# Snapshots 11 5 10 21
Node Type 1 3 4 2
Edge Type 3 3 4 3
Time Granularity Second Second Year Day
B.2.2. Dataset Processing
After the time window is partitioned, further data cleansing is necessitated. In our model and the baselines, representation vectors of nodes are derived from the initial $T$ network snapshots. However, it is not feasible to obtain representation vectors for nodes that fail to appear in the $(T+1)$-th network snapshots, rendering link prediction for these nodes impossible. Consequently, nodes that newly appear, along with their corresponding links in the $(T+1)$-th snapshot, are eliminated. Additionally, the negative links for the training set are collected from links that are absent in the initial $T$ snapshots. For the test set, negative links are sampled from those not present in the $(T+1)$-th snapshot. The node representations are derived from the first $T$ network snapshots $\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots,\mathcal{G}^{T}\}$. The links in the subsequent network snapshot $\mathcal{G}^{T+1}$ serve as the evaluation set. Moreover, 20% of the links in this evaluation set are randomly assigned as the validation set, another 20% as the positive training set, and the remaining 60% as the positive test set. Simultaneously, several non-existent negative links are randomly sampled to form negative training and test sets.
B.2.3. Metrics
We employ the widely-adopted metrics AUC and AP for evaluation. AUC represents the area under the Receiver Operating Characteristic (ROC) curve, which plots the false positive rate on the x-axis and the true positive rate on the y-axis. AP refers to the area under the Precision-Recall curve, with recall on the horizontal axis and precision on the vertical axis. Values of AUC and AP closer to 1 indicate superior model performance.
B.3. Baselines
The baselines encompass both traditional and advanced link prediction models that are highly relevant to our research. Detailed descriptions of the fundamental elements of the baseline models and our CLP are presented in Table 3.
(1) Static Homogeneous approaches:
  • SEAL (zhang2018link) extracts local graphs for the target link and learns their features to estimate the probability of that link’s existence.
  • VGNAE (ahn2021variational) integrates variational inference, GCNs, and normalization techniques to effectively learn probabilistic node embeddings from graph-structured data.
(2) Static Heterogeneous approaches:
  • Metapath2Vec (dong2017metapath2vec) leverages metapath-guided random walks combined with the skip-gram model to learn low-dimensional embeddings for nodes in heterogeneous information networks, capturing both structural and semantic relationships within the network.
  • GATNE (cen2019representation) learns node representations in attributed multiplex networks. It successfully addresses the challenges of integrating multiple types of relationships and node attributes into a unified representation.
(3) Dynamic Homogeneous approaches:
  • TGAT (XuRKKA20) presents an inductive representation learning approach for temporal graphs, combining temporal encoding and GNNs to capture the dynamic nature of such graphs, which can be generalized to new nodes and future graph snapshots.
  • TDGNN (qu2020continuous) designs a continuous-time link prediction method for dynamic graphs, leveraging temporal encodings and attention mechanisms to enhance the predictive capabilities of GNNs.
(4) Dynamic Heterogeneous approaches:
  • THAN (li2023memory) leverages memory mechanisms and transformer architectures to capture intricate temporal and structural information of temporal heterogeneous graphs.
  • THGAT (zhang2023dynamic) combines neighborhood type modeling, neighborhood information aggregation, and time encoding technique to achieve accurate node representations.
(5) Ours: